how do i solve for finding the equation for the line that passes through 0,5 that is perpendicular to the line with the equation 8x+4y=8?
Answer by mananth(16946)

how do i solve for finding the equation for the line that passes through 0,5 that is perpendicular to the line with the equation 8x+4y=8?

0,5 that is perpendicular to the line with the equation
8x+4y=8
/4
2x+y=2
y=-2x+2
m= slope = -2
...
The perpendicular line will have a slope which is negative reciprocal of the slope of the given line.
Since m1*m2=-1
so slope of the required line will be 1/2
..
line passes through (0,5)
plug the value of slope , and the point to find b the y intercept
5=(1/2)*0+b
b=5
Equation = y=(1/2)x+5